putting x^2 + y^2 =16 and x^2 - 4y^2 = 4 into y= format so it can be put in a graphing calculator
Answer by stanbon(48569)

putting x^2 + y^2 =16 and x^2 - 4y^2 = 4 into y= format so it can be put in a graphing calculator
----------
y=sqrt(16-x^2)
y=-sqrt(16-x^2)
y=(1/4)(x^2-4)
Y=(-1/4)(x^2-4)
------------
Cheers,
Stan H>